OpenStudy (anonymous):

MEDAL AND FAN! Choose the equation of the line passing through the point (-1,3) and perpendicular to y=-1/3x+7 a. y=3x-12 b. y=3x+6 c. y=3x-6 d. y=3x

OpenStudy (anonymous):

I dont want just answers please :( I need someone to explain it to me!

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

that guy will help you

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

well dude its so easy and simply -1 , 3 means that f(-1) = 3 just plug the -1 in the x in every equation and you'll get the result the equation that equals 3 is the right one y=3x-12 = - 15 y=3x+6 = 3 y=3x-6 = - 9 y=3x = - 3 so y = 3x + 6 is the right one

OpenStudy (anonymous):

:D yay! that makes sense :3! Thank you!

OpenStudy (anonymous):

happy to help you :)